OpenStudy (anonymous):

evaluate 2b to second power - 3b+4a to second power for a=-6 and b=8 ?

OpenStudy (texaschic101):

2b^2 - 3b + 4a^2 when a = -6, and b = 8 now we sub 2(8^2) - 3(8) + 4(-6^2) = 2(64) - 24 + 4(36) = 128 - 24 + 144 = 272 - 24 = 248

OpenStudy (anonymous):

awesome ty!!

OpenStudy (texaschic101):

sure thing :)